# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ 5.1 Maximum bias of the deepest scatter estimator
which says that $\varepsilon\geq 1/3.$
(ii) If $l_{p}^{(n)}\rightarrow 0,$since $\min\left(a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right),b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\right)\geq\min\left(a_{\varepsilon,n}\left(\mathbf{w}_{n}\right),b_{\varepsilon,n}\left(\mathbf{w}_{n}\right)\right),$ then either $b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)$ $\geq a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ or $a_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq b_{\varepsilon,n}\left(\mathbf{v}_{p}^{(n)}\right)\geq D\left(\Gamma_{n},P_{n}\right)$ which yields that
$\left(1-\varepsilon\right)P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{n}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}\leq l_{p}^{(n)}\right)$
$+\varepsilon P_{n}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{n}}{\left\|\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}\right\|}\leq l_{p}^{(n)}\right)$
$\geq$ $D\left(\Gamma_{n},P_{n}\right)$
Thus, since $P_{0}\left(\frac{\mathbf{w}_{n}^{t}\Gamma_{n}^{-1/2}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}XX^{t}\frac{\Gamma_{n}^{-1/2}\mathbf{w}_{r}}{\left\|\mathbf{w}^{t}\Gamma_{n}^{-1/2}\right\|}\leq l_{p}^{(n)}\right)\rightarrow 0$, we get
which says that $\varepsilon\geq 1/3.$
Given a unit vector $\mathbf{e}$ and $r>0,$ take the point mass contaminations $\delta_{r}=\delta_{r\mathbf{e}}$. If we have $P_{\varepsilon,r}=\left(1-\varepsilon\right)P_{0}+\varepsilon\delta_{r}$ and $\delta$ stands for the Kronecker delta (it is $1$ if the inequality holds, 00 otherwise), we consider
$m_{r}\left(\mathbf{u}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)P\left(-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\leq\mathbf{u}^{t}X\leq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+\varepsilon\delta\left(r^{2}\left(\mathbf{u}^{t}\mathbf{e}\right)^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ \left(1-\varepsilon\right)\left\{P\left(\mathbf{u}^{t}X\leq-\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)+P\left(\mathbf{u}^{t}X\geq\sqrt{\mathbf{u}^{t}\Gamma\mathbf{u}}\right)\right\}+\varepsilon\delta\left(r^{2}\left(\mathbf{u}^{t}\mathbf{e}\right)^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\}.$
Notation 14
Put
$h^{\mathbf{e}}\left(\mathbf{v}\right)$ $=$ $\mathbf{v}^{t}\Gamma\mathbf{v/}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}$
$A_{r}^{\mathbf{e}}$ $=$ $\left\{\mathbf{v}\in S^{p-1\text{ }}\text{and }h^{\mathbf{e}}\left(\mathbf{v}\right)>r^{2}\right\}$
and
If either $B_{r}^{\mathbf{e}}=\emptyset$ or $A_{r}^{\mathbf{e}}=\emptyset$ put $gb_{r}^{\mathbf{e}}=\infty$ and $Gb_{r}^{\mathbf{e}}=-\infty$ or $ga_{r}^{\mathbf{e}}=\infty$ and $Ga_{r}^{\mathbf{e}}=-\infty$
Lemma 15
It holds that
Proof. if $h^{\mathbf{e}}\left(\mathbf{v}\right)<r^{2},$ then $\left(1-\varepsilon\right)g\left(\mathbf{v}\right)+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\geq r^{2}\right)\geq\left(1-\varepsilon\right)gb_{r}^{\mathbf{e}}$ and $\left(1-\varepsilon\right)(1-g\left(\mathbf{v}\right))+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\leq r^{2}\right)\geq\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon.$ if $h^{\mathbf{e}}\left(\mathbf{v}\right)>r^{2},$ then $\left(1-\varepsilon\right)g\left(\mathbf{v}\right)+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\geq r^{2}\right)\geq\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon$ and $\left(1-\varepsilon\right)(1-g\left(\mathbf{v}\right))+\varepsilon\delta\left(h^{\mathbf{e}}\left(\mathbf{v}\right)\leq r^{2}\right)\geq\left(1-\varepsilon\right)\left(1-Ga_{r}^{\mathbf{e}}\right)$ and the lemma follows.
Corollary 16
Take $\Gamma$ a matrix whose eigenvector $\mathbf{v}_{1}$coincides with $\mathbf{e.}$ If $r>l_{1}^{1/2}$ consider $\mathbf{v}_{0,r}\mathbf{\in}\mathcal{S}^{p-1}$ such that $\mathbf{v}_{0,r}^{t}\left(\Gamma-r^{2}\mathbf{v}_{1}\mathbf{v}_{1}^{t}\right)\mathbf{v}_{0,r}=0.$Then
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\left(1-\varepsilon\right)(1-g(\mathbf{v}_{1})),\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\left(1-\varepsilon\right)(1-g(\mathbf{v}_{1}))+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right..$
Proof. Observe that $q_{1}=\frac{l_{1}^{1/2}}{\mathbf{v}_{1}^{t}\mathbf{e}}=l_{1}^{1/2}$ and $q_{j}=\infty,$ $j=2...,p.$
If $r<l_{1}^{1/2}$ then
@@
$\left\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{p}\right\}\in A_{r}^{\mathbf{e}}\text{ then }Ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right)\text{and }ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
@@
$\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}\notin B_{r}^{e},\mathbf{v}_{1}\in B_{r}^{e},\text{ }Gb_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right).$
Take now $\mathbf{v}_{0,r}\mathbf{\in}\mathcal{S}^{p-1}$ such that $\mathbf{v}_{0,r}^{t}\left(\Gamma-r^{2}\mathbf{v}_{1}\mathbf{v}_{1}^{t}\right)\mathbf{v}_{0,r}=0.$ Since $g\left(\mathbf{v}_{2}\right)\leq g\left(\mathbf{v}_{0,r}\right)<g\left(\mathbf{v}_{1}\right),$
$gb_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{0,r}\right).$
Observe that $\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}\in A_{r}^{\mathbf{e}}$ and $\mathbf{v}_{1}\notin A_{r}^{\mathbf{e}}.$ Then
@@
$Ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{0,r}\right)\text{ and }ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
Since
$D\left(\Gamma,P_{\varepsilon,r}\right)$ $=$
$\min\left\{\left(1-\varepsilon\right)\left(1-Gb_{r}^{\mathbf{e}}\right)+\varepsilon,\left(1-\varepsilon\right)ga_{r}^{\mathbf{e}}+\varepsilon,\left(1-\varepsilon\right)gb_{r}^{\mathbf{e}},\left(1-\varepsilon)\left(1-Ga_{r}^{\mathbf{e}}\right)\right)\right\},$
then we have that
@@
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\\ \left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\end{array}\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right..$ (17)
The following theorem restricted to matrices whose eigenvector associated with the largest eigenvalue shares the same direction than the contamination gives us which should be the deepest estimator in this class.
Lemma 17
Given $\Gamma=\sum l_{j}\mathbf{v}_{j}\mathbf{v}_{j}^{t}$, $l_{1}\geq...\geq l_{p},$ $\mathbf{v}_{i}^{t}\mathbf{v}_{j}=\delta_{ij}$ and $\mathbf{v}_{1}=\mathbf{e.}$ Then the deepest matrices in this class are given by
with $\beta=\left[\Phi^{-1}\left(\frac{3-4\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},$ $l_{1}=\left[\Phi^{-1}\left(\frac{3-\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},l_{p}=\left[\Phi^{-1}\left(\frac{3-5\varepsilon}{4\left(1-\varepsilon\right)}\right)\right]^{2},l_{p}\leq l_{p-1}\leq\dots\leq l_{2}<\Phi^{-1}\left(\frac{3}{4}\right)$ and $D\left(\Gamma,P_{\varepsilon,r}\right)=\left(1-\varepsilon\right)/2.$
Proof. If $r<l_{1}^{1/2}$ then the deepest matrix should verify
@@
$g\left(\mathbf{v}_{p}\right)+g\left(\mathbf{v}_{1}\right)$ $=$ $\frac{1-2\varepsilon}{1-\varepsilon}$
To increase the depth, $g\left(\mathbf{v}_{p}\right)$ should increase and $g\left(\mathbf{v}_{1}\right)$ decrease, and then they should reach $1/2$ This implies that
@@
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right),\\ \left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)\right)\end{array}\right\}.$
with
$g(\mathbf{v}_{p})\leq...\leq g(\mathbf{v}_{2})<g\left(\mathbf{v}_{0,r}\right)<g\left(\mathbf{v}_{1}\right).$
Given $\Gamma$ we have for any $l_{p}\leq$ $s\leq l_{1},$ $l_{p}/r^{2}\leq$ $s/r^{2}\leq l_{1}/r^{2}<1$ and the equation $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=s=r|\mathbf{v}_{0,r}^{t}\mathbf{v}_{1}|$ has a solution $\mathbf{v}_{0,r}.$ Geometrically, $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=s=r|\mathbf{v}_{0,r}^{t}\mathbf{v}_{1}|$ means that $\mathbf{v}_{0,r}$ belongs to the intersection of an ellipsoid and a cone around $\mathbf{v}_{1}.$ $\mathbf{v}_{0,r}$ is a function of $r,$ and $\mathbf{v}_{0,r}\underset{r\rightarrow\infty}{\rightarrow}\mathbf{v}_{1}.$The best estimator should be selected so that $\left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right)=(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right)$ which entails that $g\left(\mathbf{v}_{0,r}\right)=1/2$ or $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=\mathbf{\Phi}^{-1}\left(3/4\right).$   Then we take matrices $\Gamma$ such that $\Phi^{-1}\left(3/4\right)\leq l_{1}^{1/2}$ (in special $s=\Phi^{-1}\left(3/4\right)=0.6745).$ All the four quantities involved in $D\left(\Gamma,P_{\varepsilon,r}\right)$ must be equal in order to get the best estimator In this case we get that
@@
$\left(1-g_{1}\right)+\frac{\varepsilon}{1-\varepsilon}$ $=$ 12 ; $\frac{1}{2}\text{ \ };\text{ \ }g_{p}+\frac{\varepsilon}{1-\varepsilon}=\frac{1}{2}$
$g_{1}$ $=$ $\frac{1}{2}+\frac{\varepsilon}{1-\varepsilon}\text{ \ \ ; \ \ }g_{p}=\frac{1}{2}-\frac{\varepsilon}{1-\varepsilon}$
$g_{1}$ $=$ $\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\text{ \ \ ; \ \ }g_{p}=\frac{1-2\varepsilon}{2\left(1-\varepsilon\right)}.$

### v2

@@ 7.1 Maximum bias of the deepest scatter estimator
which says that $\varepsilon\geq 1/3.\quad\square$
[…]
Lemma 7.2
The critical points of $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)$ (respectively $1-\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right))$ subject to $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\geq 0$ (idem $\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\leq 0)$ are either $\mathbf{v}_{p},$ $\mathbf{v}_{1}$ or occurs at $F_{r}^{\mathbf{e}}.$
Proof. Take the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\lambda\left(\mathbf{v}^{t}\Gamma\mathbf{v-}r^{2}\left(\mathbf{v}^{t}\mathbf{e}\right)^{2}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right),$ Karush, Kuhn, Tucker condition says that if the minimum local occurs in $\mathbf{\tilde{v},}$ then, either $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}$ == 0 or we have to consider the Lagrangian $\Phi\left(\sqrt{\mathbf{v}^{t}\Gamma\mathbf{v}}\right)+\gamma\left(\mathbf{v}^{t}\mathbf{v-}1\right).$ In the last case, it was shown that the critical points are the eigenvectors of $\Gamma.$ Suppose that $\mathbf{v}_{j_{0}\text{,}}$ $j_{0}<p,$is the eigenvector corresponding to the minimum $l_{j_{0}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$, by the continuity of the functions we could find another $\mathbf{\bar{v}}$ such that $\mathbf{v}_{j_{0}\text{ }}^{t}\Gamma\mathbf{v}_{j_{0}\text{ }}>\mathbf{\bar{v}}_{\text{ }}^{t}\Gamma\mathbf{\bar{v}}>r^{2}\left(\mathbf{\bar{v}}^{t}\mathbf{e}\right)^{2}>r^{2}\left(\mathbf{v}_{j_{0}\text{ }}^{t}\mathbf{e}\right)^{2}$ which is a contradiction. Therefore the critical points occur at $\mathbf{\tilde{v}=v}_{p}$ $\left(\mathbf{v}_{1}\right)$ or $\mathbf{\tilde{v}}^{t}\Gamma\mathbf{\tilde{v}-}r^{2}\left(\mathbf{\tilde{v}}^{t}\mathbf{e}\right)^{2}=0.$
Proof of Corollary 4.6 Set $q_{j}=\frac{l_{j}^{1/2}}{\mathbf{v}_{j}^{t}\mathbf{e}},$ then $q_{1}=l_{1}^{1/2}$ and $q_{j}=\infty,$ j=2,p.j=2...,p. If $r<l_{1}^{1/2}$ then
@@
$\left\{\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{p}\right\}\subseteq A_{r}^{\mathbf{e}}\text{ then }Ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right)\text{ and }ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
@@
$\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}$ $\subseteq$ $\left(B_{r}^{\mathbf{e}}\right)^{c},\mathbf{v}_{1}\in B_{r}^{\mathbf{e}},\text{ }Gb_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{1}\right),$
$\left\{\mathbf{v}_{2},...,\mathbf{v}_{p}\right\}$ $\subseteq$ $A_{r}^{\mathbf{e}},\text{ }\mathbf{v}_{1}\notin A_{r}^{\mathbf{e}},ga_{r}^{\mathbf{e}}=g\left(\mathbf{v}_{p}\right).$
We need to determine $gb_{r}^{\mathbf{e}}$ and $Ga_{r}^{\mathbf{e}}.$ Then the function $g\left(\mathbf{v}\right)$ has a minimum in $B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}$ (a compact set) and the minimum has to occur in $F_{r}^{\mathbf{e}}.$ Likewise, the function $g\left(\mathbf{v}\right)$ has a maximum in $A_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}$ which must occur at $F_{r}^{\mathbf{e}}.$ In both cases, Lemma 7.2 supports that statement. We use spherical coordinates,
$\mathbf{v}^{t}\mathbf{v}_{1}$ $=$ $\cos\eta_{0},\text{ }$
$\mathbf{v}^{t}\mathbf{v}_{2}$ $=$ $\sin\eta_{0}\cos\eta_{1},$
$\mathbf{v}^{t}\mathbf{v}_{3}$ $=$ $\sin\eta_{0}\sin\eta_{1}\cos\eta_{2},\text{ }0\leq\eta_{0},..\eta_{p-3}\leq\pi$
$\vdots$
$\mathbf{v}^{t}\mathbf{v}_{p-1}$ $=$ $\sin\eta_{0}\sin\eta_{1}\sin\eta_{2}\dots\cos\eta_{p-2},0\leq\eta_{p-2}\leq 2\pi$
$\mathbf{v}^{t}\mathbf{v}_{p}$ $=$ $\sin\eta_{0}\sin\eta_{1}\sin\eta_{2}\dots\sin\eta_{p-2},$
($\eta_{j}$ is the angle between $\mathbf{v}_{j+1}$ and $\mathbf{v,}$ $j=0,...,p-1\mathbf{).}$ Then
$\mathbf{v}^{t}\Gamma\mathbf{v=}r^{2}\left(\mathbf{v}^{t}\mathbf{v}_{1}\right)^{2}$
is equivalent to
$l_{1}\cos^{2}\eta_{0}+\sin^{2}\eta_{0}\left[l_{2}\cos^{2}\eta_{1}+l_{3}\sin^{2}\eta_{1}\cos^{2}\eta_{2}+\dots+l_{p}\sin^{2}\eta_{1}\sin^{2}\eta_{2}\dots\sin^{2}\eta_{p-2}\right]=r^{2}\cos^{2}\eta_{0}$
$tg^{2}\eta_{0}=\frac{r^{2}-l_{1}}{(l_{2}\cos^{2}\eta_{1}+\sin^{2}\eta_{1}(\dots(l_{p-2}\cos^{2}\eta_{p-3}+\sin^{2}\eta_{p-3}\left(l_{p-1}\cos^{2}\eta_{p-2}+l_{p}\sin^{2}\eta_{p-2}\right))...)}$
Observe that the denominator are nested convex combinations. Thus, we can ensure that
$\frac{r^{2}-l_{1}}{l_{2}}$ $\leq$ $tg^{2}\eta_{0}\leq\frac{r^{2}-l_{1}}{l_{p}}$
$\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}$ $\leq$ $\left|tg\left(\eta_{0}\right)\right|\leq\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}$
$\text{either }\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}$ $\leq$ $tg\left(\eta_{0}\right)\leq\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\text{ or }-\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\leq tg\left(\eta_{0}\right)\leq-\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}$
$\text{either }0$ $\leq$ $\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\leq\eta_{0}\leq\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\leq\pi/2\text{ }$
$\text{or }\pi/2$ $\leq$ $\arctan\left(-\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\leq\eta_{0}\leq\arctan\left(-\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\leq\pi$
Then, by similarity of triangles we can say that, for either $l=l_{2}$ or $l_{p},$
$\frac{\sqrt{1+\frac{r^{2}-l_{1}}{l}}}{1}=\frac{1}{\cos\left(\theta\right)},\text{ \ \ \ }\frac{\sqrt{1+\frac{r^{2}-l_{1}}{l}}}{1}=\frac{\sqrt{\frac{r^{2}-l_{1}}{l}}}{\sin\left(\theta\right)}$
Therefore, we obtain that
@@
$\max_{A_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$ $\mathbf{=}$ $\max_{F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v=}r^{2}\cos^{2}\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\right]=r^{2}\frac{l_{2}}{r^{2}+l_{2}-l_{1}}\text{ }$
$\mathbf{v}_{M,r}$ $=$ $\arg\max_{F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$
$\mathbf{=}$ $\left(\cos\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right),\sin\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right),0,\dots,0)\right)$
$=$ $\left(\sqrt{\frac{l_{2}}{r^{2}+l_{2}-l_{1}}},\sqrt{\frac{r^{2}-l_{1}}{r^{2}+l_{2}-l_{1}}},0,\dots,0\right)$
$\min_{B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$ $\mathbf{=}$ $r^{2}\cos^{2}\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\right]=r^{2}\frac{l_{p}}{r^{2}+l_{p}-l_{1}}$
$\mathbf{v}_{m,r}$ $=$ $\arg\min_{B_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}}}\mathbf{v}^{t}\Gamma\mathbf{v}$
$\mathbf{=}$ $\left(\cos\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right),0,\dots,0,\sin\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\right)$
$=$ $\left(\sqrt{\frac{l_{p}}{r^{2}+l_{p}-l_{1}}},0,\dots,0,\sqrt{\frac{r^{2}-l_{1}}{r^{2}+l_{p}-l_{1}}}\right)$
which means that
$gb_{r}^{\mathbf{e}}$ $=$ $g\left(\mathbf{v}_{m,r}\right)=2\Phi\left(r\cos\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{p}}}\right)\right]\right)-1,$
$Ga_{r}^{\mathbf{e}}$ $=$ $g\left(\mathbf{v}_{M,r}\right)=2\Phi\left(r\cos\left[\arctan\left(\sqrt{\frac{r^{2}-l_{1}}{l_{2}}}\right)\right]\right)-1.$
Since $\mathbf{v}_{M,r}$ is a maximum over $A_{r}^{\mathbf{e}}\cup F_{r}^{\mathbf{e}},$ $g\left(\mathbf{v}_{2}\right)\leq g\left(\mathbf{v}_{M,r}\right)\leq g\left(\mathbf{v}_{1}\right).$ Then, we have that
@@
$D\left(\Gamma,P_{\varepsilon,r}\right)=\left\{\begin{array}[]{cc}\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)g(\mathbf{v}_{p})+\varepsilon,\\ \left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)\end{array}\right\}&\text{if }r\leq l_{1}^{1/2}\\ \min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{m,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{M,r}\right)\right)\right)\end{array}\right\}&\text{if }r>l_{1}^{1/2}\end{array}\right.$. (30)
[…]
Proof of Lemma 4.7 If $r<l_{1}^{1/2}$ then the deepest matrix should verify
@@
$g\left(\mathbf{v}_{p}\right)+g\left(\mathbf{v}_{1}\right)$ $=$ $\frac{1-2\varepsilon}{1-\varepsilon}.$
To increase the depth, $g\left(\mathbf{v}_{p}\right)$ should increase and $g\left(\mathbf{v}_{1}\right)$ should decrease, and then they would reach
$2g\left(\mathbf{v}_{p}\right)=2g\left(\mathbf{v}_{1}\right)=\frac{1-2\varepsilon}{1-\varepsilon}.$
This implies that
@@
$D\left(\Gamma,P_{\varepsilon,r}\right)=\min\left\{\begin{array}[]{c}\left(1-\varepsilon\right)\left(1-g\left(\mathbf{v}_{1}\right)\right)+\varepsilon,\left(1-\varepsilon\right)g\left(\mathbf{v}_{p}\right)+\varepsilon,\\ \left(1-\varepsilon\right)g\left(\mathbf{v}_{m,r}\right),\left(1-\varepsilon)\left(1-g\left(\mathbf{v}_{M,r}\right)\right)\right)\end{array}\right\}.$
with $g(\mathbf{v}_{p})\leq...\leq g(\mathbf{v}_{2})\leq g\left(\mathbf{v}_{M,r}\right)\leq g\left(\mathbf{v}_{1}\right).$ The best estimator should be selected so that $g\left(\mathbf{v}_{m,r}\right)=g\left(\mathbf{v}_{M,r}\right)$ (by making $l_{2}=\dots=l_{p},$ $g\left(\mathbf{v}_{m,r}\right)$ increases) and $\mathbf{v}_{m,r}=\mathbf{v}_{M,r}.$ Call $\mathbf{v}_{0,r}=\mathbf{v}_{m,r}$ and $\left(1-\varepsilon\right)g\left(\mathbf{v}_{0,r}\right)=(1-\varepsilon)\left(1-g\left(\mathbf{v}_{0,r}\right)\right),$ which entails that $g\left(\mathbf{v}_{0,r}\right)=1/2$ or $\sqrt{\mathbf{v}_{0,r}^{t}\Gamma\mathbf{v}_{0,r}}=\mathbf{\Phi}^{-1}\left(3/4\right).$  Then we take matrices $\Gamma$ such that $\Phi^{-1}\left(3/4\right)=0.6745\leq l_{1}^{1/2}$ . All the four quantities involved in $D\left(\Gamma,P_{\varepsilon,r}\right)$ must be equal in order to get the best estimator In this case we get that
@@
$\left(1-g_{1}\right)+\frac{\varepsilon}{1-\varepsilon}$ $=$ 12 ; $\frac{1}{2}\text{ \ };\text{ \ }g_{p}+\frac{\varepsilon}{1-\varepsilon}$$\frac${1}{2}
[…]
$g_{1}$ $=$ $\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\text{ \ \ ; \ \ }g_{p}=\frac{1-2\varepsilon}{2\left(1-\varepsilon\right)},$